HUBBLE SPACE TELESCOPE DAILY REPORT #856 PERIOD COVERED: 0800 EST Daily Status Report as of 091/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: Safemode was officially ended at 091/000Z, and SMS 091 was successfully initiated. 1.1 Completed XSAFE (Safemode Recovery of the HST Science instruments - FOC) The XSAFE Safemode recovery was successful. No problems were reported. 1.1 Completed WQSAF (Safemode Recovery of the HST Science instruments - WFPC) The WQSAF Safemode recovery was successful. No problems were reported. 1.2 Completed YSAFE (Safemode Recovery of the HST Science instruments - FOS) The YSAFE Safemode recovery was successful. No problems were reported. 1.3 completed VSAFE (Safemode Recovery of the HST Science instruments - HSP) The VSAFE Safemode recovery was successful. No problems were reported. 1.4 Completed FOS/RD 2563 (SINS: The Super Nova Intensive Study: Cycle 1 Observations) The Faint Object Spectrograph was used to make observations of the supernova LMC-SN-1987A which is located in the Large Magellanic Cloud. The observations were completed as scheduled, and no problems were reported. 1.5 Completed PC 4347 (WF/PC Cycle 2 Calibration: Flood retention Test) The Planetary Camera was used to make observations looking for changes in the quantum efficiency hysteresis since the last decontamination operation. The observations were completed as scheduled, and no problems were reported. 1.6 Initiated FOS/RD 4123 (FOS Spectral Flat Fields Monitor, Revised) The Faint Object Spectrograph (Red Side) was used to initiated observations of object WD0501+527. The observations are currently under way, and no problems have been reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 5 Successful : 5 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 7 Successful : 7 2.3 Significant HSTARs: None 2.4 Operations Note: Safemode was officially ended at 091/000Z, and SMS 091 was successfully initiated. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/RD 4123 (FOS Spectral Flat Fields Monitor, Revised) 3.2 Complete PC 3603 (Snapshot of Protoplanetary Nebulae) 3.3 Complete HRS/2 4806 (Second Epoch High Resolution Ultraviolet Spectroscopy of Nova Cygni 19) 3.4 Complete FOS/BL 2563 (SINS: The Super Nova Intensive Study: Cycle 1 Observations) 3.5 Uplink Updated Ephemeris Tables, Per MOSES 2871 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-